PUBLICATIONS

1. Myasnikov A. On the approximability of groups of outer automorphisms of free groups of finite rank. Algebra and Logic, 1981, v.20, 3, p.291-299.

2. Myasnikov A., Remeslennikov V. Isomorphism and elementary properties of powered nilpotent groups. Proc.Math.Institute, v.2, Nauka, Novosibirsk, 1981, p.56-87

3. Myasnikov A., Remeslennikov V. Elementary properties of powered nilpotent groups. Doklady Akademii Nayuk, 1981,v. 258, 5, p.1056-1059.

4. Myasnikov A., Remeslennikov V., Definability of Mal’cev bases and elementary properties of finite dimensional algebras.I, Siberian Jour. Math., 1982, v.23, 5, p.152-167.

5. Myasnikov A., Remeslennikov V. Definability of Mal‘cev bases and elementary theories of finite dimensional algebras.II, Siberian Jour. Math., 1983, v.24, 2, p.97-113.

6. Myasnikov A. Finite-dimensional algebras and k-groups of finite rank. Contemp Math., 1984, v.33, p.436-453.

7. Myasnikov A. Generalized Nielsen transformations and trivial groups. Math. Notes, 1984, v.5, 4, p.491-495.

8. Myasnikov A. Stable torsion-free rings and nilpotent groups. In: VII Soviet Union conf. on math. logic, Novosibirsk, 1984.

9. Myasnikov A., Remeslennikov V. Elementary group equivalence with the integral length function. Illinois J.Math., 20,1986, v.30,2.

10. Myasnikov A. Free operator groups with length function and finitely generated group of operators. In: Computable invariants of algebraic structures.Novosibirsk, 1987, p.87-97.

11. Myasnikov A. Abstract isomorphisms and elementary theories of finite dimensional algebras and unipotent groups. Doklady Akademii Nauk USSR, 1987, v.297, 2, p.290-293.

12. Myasnikov A., Remeslennikov V. Recursive p-adic numbers and elementary theories of finitely generated pro-p-groups. Izvest. Akademii Nauk USSR, ser.math 1988, v.30, 3.

13. Myasnikov A. Structure of models and criteria of decidability of complete theories of finite-dimensional algebras. Izvest. Akademii Nayk USSR, ser. math., 1989, v.53,2, p.379-397.

14. Myasnikov A. Modules over local ring. Siberian Jour. Math., 1989, v.30,3, p.72-83.

15. Myasnikov A. Remeslennikov V. Model-theoretical problems in group theory. In: Problems of Algebra, 1989, Minsk, v.4.

16. Myasnikov A. Invariants of bilinear mappings. Siberian Jour. Math., 1990, v.31,1, p.104-115.

17. Myasnikov A. Model theory of bilinear mappings. Siberian Jour. Math., 1990, v.31,3, p.94-108.

18. Myasnikov A.,Remeslennikov V. Logic of finite types in group theory. Bull. Siberian Math Soc., Novosibirsk, 1990, p.18-25.

19. Myasnikov A., Remeslennikov V. Weak second order logic in group theory. Contemp. Math., 1992(part 2), v.131, p.273-277.

20. Myasnikov A., Remeslennikov V. Admissible sets in group theory. Algebra i logic, 1992,v.3.

21. Myasnikov A.G. Model–theoretic problems of algebra. Dissertation, VINITI, Moscow, 1992.

22. Myasnikov A., Belyaev V. Recursively enumerable degrees over the reals. In Proc. of Intern. Algebraic Conf., Krasnoyarsk, 1993.

23. Ashaev I., Belyaev V., Myasnikov A. Approaches to the theory of computability. Algebra i Logic, 1994, v.4.

24. Myasnikov A.G., Remeslennikov V.N. Exponential groups I: foundations of the theory and tensor completion. Siberian Math. J., 1994, v.35, N. 5, p.1106-1118.

25. O.Kharlampovich, A. Myasnikov and V. Remeslennikov. Logical languages and axioms for groups with a length function. Proc. Inst. Appl. Math. Russian Acad. Sci., 1995, 20, p.1-8.

26. D. Gildenhuys, O. Kharlampovich and A. Myasnikov. CSA groups and separated free constructions. Bull. Austr. Math. Soc., 1995, v. 52, 1, p.63-84.

27. Myasnikov A.G., Remeslennikov V.N. Length functions on free exponential groups. Proc. of Intern. Conference Groups in Analysis and Geometry, Omsk, 1995, p.59-61.

28. A.G.Myasnikov, V.N.Remeslennikov. Length functions on free exponential groups. Proc. N 26. IITPM SO RAN, Omsk, 1996, 1-34.

29. A. Myasnikov and V. Remeslennikov. Exponential groups 2: extension of centralizers and tensor completion of CSA-groups. Intern. Journal of Algebra and Computation, Vol.6, No.6 (1996), p.687-711.

30. A.M. Gaglione, A.G. Myasnikov, V.N. Remeslennikov, D. Spellman. Formal Power Series Representations of Free Exponential Groups. Communications in Algebra, 1997, v. 25(2), p.631-648.

31. O. Kharlampovich and A. Myasnikov. Hyperbolic groups and free constructions. Transactions of Math., 350(1998), 2, pp.571-613.

32. O.Kharlampovich, A. Myasnikov. Equations in a free Q-group. Transactions of Math., 350(1998), 3, pp. 947-974.

33. O.Kharlampovich, A.Myasnikov. Irreducible affine varieties over a free group. I: Irreducibility of quadratic equations and Nullstellensatz. J. of Algebra, 1998, v. 200, n. 2, p.472-516.

34. O.Kharlampovich, A.Myasnikov. Irreducible affine varieties over a free group. II: Systems in triangular quasi-quadratic form and description of residually free groups. J. of Algebra, 1998,v. 200, n. 2, p. 517-570.

35. A.Gaglione, B.Fine, A.Myasnikov, G.Rosenberger, D.Spellman. A classification of fully residually free groups of rank three or less. Journal of Algebra, 1998, v. 200, n. 2, p. 571-605.

36. O.Kharlampovich, A.Myasnikov. Description of Fully Residually Free Groups and irreducible affine varieties over a free group. CRM Proceedings and Lecture Notes, 1998, v. 17, p. 71-80.

37. G.Baumslag, A.Myasnikov, V.Remeslennikov. Algebraic geometry over groups I. Algebraic sets and ideal theory. Journal of Algebra, 1999, v.219, pp.16-79.

38. O.Kharlampovich, E. Lyutikova, A. Myasnikov. Equations over Qcompletions of hyperbolic groups. Transactions of Math., 1999.

39. G.Baumslag, A.G.Myasnikov, V. Shpilrain, Open problems in combinatorial group theory. In: Groups, Languages and Geometry. Amer. Math. Soc., Contemporary Math. 250, 1999, pp.1-27.

40. G.Baumslag, A.Myasnikov, V.Remeslennikov. Malnormality is decidable in free groups. Intern. Journal of Algebra and Computation, v.9, 6, 1999.

41. K.Bencsath, B.Fine, A.Gaglione, A.Myasnikov, G.Rosenberger, D.Spellman. Aspects of the theory of free groups. Algorithmic problems in groups and semigroups, Birkhauser, Trends in Math., Boston, Basel, Berlin, pp.35-50.

42. G.Baumslag, A.Myasnikov, V.Remeslennikov. Algebraic geometry over groups. Algorithmic problems in groups and semigroups, Birkhauser, Trends in Math., Boston, Basel, Berlin, 2000, pp.51-90.

43. S.Lioutikov and A.Myasnikov. Centroids of groups Journal of Group Theory, v.3, iss.2 (2000), pp. 177-197.

44. A.Myasnikov, V.Remeslennikov. Algebraic geometry 2: logical foundations. Journal of Algebra 234 (2000), p. 225-276.

45. G.Baumslag, A.Myasnikov, V.Remeslennikov. Discriminating and codiscriminating groups. J. of Group Theory, v. 3, no. 4 (2000), p. 467479 .

46. A.D.Miasnikov and A.G.Myasnikov. Balanced presentations of the trivial group on two generators and the Andrews-Curtis conjecture, Groups and Computation III, Walter de Gruyter, Berlin New York, 2001, p.257-263.

47. O.Kharlampovich, A. Myasnikov. Implicit function theorem over free groups and genus problem. AMS/IP Studies in Advanced Mathematics, v.24, 2001, pp.77-83.

48. B.Fine, A.M.Gaglione, A.Myasnikov, D.Spellman. Discriminating groups, J. of Group Theory 4 (2001), pp.463-474.

49. I.Kapovich and A.Myasnikov Stallings foldings and subgroups of free groups. Journal of Algebra, 248 (2002), pp.608–668.

50. A.D. Miasnikov, A. G. Myasnikov, V. Shpilrain On the Andrews-Curtis equivalence, In: Combinatorial and Geometric Group Theory. Amer. Math. Soc., Contemp. Math., 296, (2002), pp.183-198.

51. G. Baumslag, A.G. Myasnikov, V. Shpilrain, Open problems in combinatorial group theory. Second edition. In: Combinatorial and Geometric Group Theory. Amer. Math. Soc., Contemporary Math. 296 (2002), pp.1-38.

52. A. Borovik, A. Myasnikov, V. Shpilrain Measuring sets in infinite groups. In: Computational and Statistical Group Theory. Amer. Math. Soc., Contemporary Math. 298 (2002), pp.21-42.

53. G. Baumslag, A. Myasnikov, V. Remeslennikov. Discriminating completions of hyperbolic groups, Geometriae Dedicata (2002), p.1-29.

54. I.Kapovich, A.Myasnikov, P.Schupp, V.Shpilrain Generic-case complexity and decision problems in group theory J. of Algebra, 264 (2003), pp.665-694.

55. A.Myasnikov, V.Shpilrain Automorphic orbits in free groups, J. of Algebra Volume, 269 (2003),Issue 1, pp. 18-27.

56. A. V. Borovik, E. I. Khukhro and A. G. Myasnikov The Andrews-Curtis Conjecture and Black Box Groups, Int. J. of Algebra and Computation, v.13, n.4 (2003), p.415-436.

57. A. Borovik, A. Myasnikov, V. Remeslennikov Multiplicative measures on free groups, International Journal of Algebra and Computation, v.13, n.6 (2003), p.705-731.

58. Ilya Kapovich, Alexei Myasnikov, Richard Weidmann Foldings, graphs of groups and the membership problem, Int. J. of Algebra and Computation, 2003.

59. B. Fine, A. Gaglione, A. Myasnikov, D. Spellman Elementary theory of free groups, Proceedings of Int. Conf on Combinatorial group theory, Oxford, 2001.

60. R.Gilman and A.Myasnikov One Variable Equations in Free Groups via Context Free Languages, Contemporary Mathematics, v. 349, 2004, p. 83-88.

61. A.Myasnikov and P. Shumyatsky Discriminating groups and c-dimension, Journal of Group Theory, Volume 7, Issue 1 (2004), p. 135142.

62. Alexei D. Miasnikov, Alexei G. Myasnikov Whitehead method and Genetic Algorithms, Contemporary Mathematics, Contemporary Mathematics, v. 349, 2004, p. 89-114.

63. Robert M. Haralick, Alexei D. Miasnikov, Alexei G. Myasnikov Pattern Recognition Approaches to Solving Combinatorial Problems in Free Groups, Contemporary Mathematics, v. 349, 2004, p. 197-213.

64. O. Kharlampovich, A.Myasnikov, V. Remeslennikov, D. Serbin Subgroups of fully residually free groups: algorithmic problems, Contemporary Mathematics, 2004.

65. I.Kapovich, A.Myasnikov, P.Schupp, V.Shpilrain Average-case complexity for the word and membership problems in group theory. Adv. Math.

66. Belyaev V., Myasnikov A. Theorems on hierarchy in generalized recursion theory.

67. A.D. Miasnikov, A.G. Myasnikov and R.M Haralick Heuristics for Whitehead Minimization Problem.

68. O.Kharlampovich, A. Myasnikov. Elementary theory of free groups. The Tarski problem.

69. T.Jitsukawa, B.Khan, A.Myasnikov On the Hanna Newmann Conjecture. Preprint.

70. A. Kvaschuk, A. Myasnikov One variable equations over hyperbolic groups. Preprint.

71. O.Kharlampovich, A.Myasnikov, V.Remeslennikov, D.Serbin G-constructible groups. Submitted.

72. I.Bumagin, O.Kharlampovich, A.Myasnikov Isomorphism problem for finitely generated fully residually free groups. Submitted.

73. Martin Kreuzer, Alexei Myasnikov, Gerhard Rosenberger, Alexander Ushakov Quotient Tests and Gröbner Bases . Preprint.

74. Alexandre V. Borovik and Alexei G. Myasnikov Quotient tests and random walks in computational group theory . Submitted.

75. Joel David Hamkins, Alexei Myasnikov The halting problem is almost always decidable. Preprint.

76. Alexei Myasnikov, Vladimir Shpilrain, Alexander Ushakov A practical attack on some braid group based cryptographic protocols. Submnitted.

77. O.Kharlampovich, A. Myasnikov. Implicit function theorems over free groups. To appear in Journal of Algebra. To appear.

78. O. Kharlampovich, A. Myasnikov, Algebraic geometry over free groups: lifting solutions into generic points, to appear in Contemporary Mathematics.

79. B.Fine, A.M.Gaglione, A.Myasnikov, D.Spellman. Groups Whose Universal Theory is Axiomatizable by Quasi-Identities, to appear in J. of Group Theory.

80. I.Kapovich, A.Myasnikov, P.Schupp, V.Shpilrain Average-case complexity for the word and membership problems in group theory. Adv. Math.

81. A. Kvaschuk, A. Myasnikov, Big powers and free constructions, Submitted.

82. A. V. Borovik, A. G. Myasnikov and V. N. Remeslennikov, Conjugacy problem in HNN-extensions of groups, to appear in Contemporary Mathematics.

83. A. V. Borovik, A. G. Myasnikov and V. N. Remeslennikov, Complexity of the conjugacy problem in amalgamated products of free groups, to appear in Contemporary Mathematics.

84. A.Myasnikov, V.Remeslennikov, D.Serbin, Regular free length functions on Lyndon’s free Z[t]-group F^{Z[t]}, to appear in Contemporary Mathematics.

85. O.Kharlampovich and A.Myasnikov Effective JSJ decompositions, to appear in Contemporary Mathematics.

86. Robert M. Haralick, Alexei D. Miasnikov, Alexei G. Myasnikov, Pattern Recognition and Minimal Words in Free Groups of Rank 2. Journal of Group Theory.

87. A.Borovik, A. Lubotzky, A.Myasnikov Finitary Andrews-Curtis conjecture. To appear.

88. A. Kvaschuk, A. Myasnikov, V.Remeslennikov Algebraic geometry over groups III: Elements of model theory. To appear in Journal of Algebra.

Articles in preparation

1. A. D. Miasnikov, A. G. Myasnikov, A. Ushakov, Andrews-Curtis Conjecture and presentations of length 14.